OpenStudy (jameshorton):

Which points are 56 units from the point located at 13 on a number line? A. –69 and 69 B. –69 and 43 C. –43 and 69 D. 43 and 69

OpenStudy (jameshorton):

@peachpi

OpenStudy (anonymous):

13 + 56 = 13 - 56 =

OpenStudy (jameshorton):

is it A

OpenStudy (anonymous):

no see above

OpenStudy (jameshorton):

it -43

OpenStudy (anonymous):

yes

OpenStudy (jameshorton):

c

OpenStudy (anonymous):

yes

OpenStudy (jameshorton):

Which integer is closest to 0 on the number line? A. –7 B. –2 C. 1 D. 3

OpenStudy (anonymous):

1 is closest to 0.

OpenStudy (anonymous):

do understand this stuff at all?

OpenStudy (jameshorton):

kinda

OpenStudy (jameshorton):

my brother wont help me

OpenStudy (anonymous):

watch a few of the videos on this site, they're pretty short and to the point, and helpfull especially if you learn things visually https://www.khanacademy.org/math/cc-seventh-grade-math/cc-7th-negative-numbers-add-and-subtract